• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 16:00:11 on September23,2009

    Entry number 291178

    keyword=Day Shift Summary (23 Sep 2009)

    Day Shift Summary (23 September 2009)

    Shift crew: Diana Parno (SL), Paul Reimer (TO), Stanley Kowalski (3rd)

    Began shift in production running at 99 uA, taking data for Slug 12 (IHWP IN), Run 12597.

    07:55 - Lost communication with LHRS magnet control IOC (iocha14). Rebooted IOC. Cycled magnets out of an abundance of caution (it turns out thi s was not necessary).

    08:30 - LHRS magnet momentum values reached their setpoint of 3.14159 GeV/c.

    08:45 - We realize CODA has not been collecting any events in Run 12598. Fixed with kcoda, startcoda.

    09:07 - Beam down due to HV trip on RF cavity. Switched to pedestal runs

    09:44 - Beam is back.

    12:18 - Beamline HV crate trips.

    14:15 - Turns out Run 12606 crashed about halfway through without anyone realizing.

    14:26 - Beamline HV crate tripped again.

    15:16 - Compton HV tripped.

    15:19 - Compton gamma rate (from Saclay DAQ) stopped updating. No complaints from CODA. I stopped run and then started it again.

    Finished shift in production running at 97.8 uA (Run 12609).

    PRODUCTION RUNS
    Run #....Good Pairs
    12597......33.0k
    12600.......5.7k
    12602......41.0k
    12603......32.6k
    12604......26.7k
    12605......32.2k
    12606......15.9k
    12608......43.3k

    PEDESTAL RUNS
    Run #
    125601

    JUNK RUNS
    Run #.
    12598
    12599
    12607